# Beyond Linearity and Time-homogeneity: Relational Hyper Event Models with Time-Varying Non-Linear Effects

### v2

@@ Abstract
List of Abbreviations: REM: Relational Event Model RHEM: Relational Hyper Event Model GAM: Generalized Additive Model TVE: Time-Varying Effect NLE: Non-Linear Effect LE: Linear Effect TVNLE: Joint Time-Varying Non-Linear Effect TPRS: Thin Plate Regression Spline TPS: Thin Plate Spline MLE: Maximum Likelihood Estimator ECDF: Empirical Cumulative Distribution Function LogLik: Log-Likelihood
@@ 1 Introduction
Activities such as sending e-mails (Perry and Wolfe, 2013; Boschi and Wit, 2024), executing financial transactions (Bianchi and Lomi, 2023), attending meetings (Lerner et al., 2021), citing academic references (Filippi-Mazzola and Wit, 2024b; Lerner et al., 2025), co-participating in criminal activities (Bright et al., 2024), and collaborating in cultural production (Burgdorf et al., 2024) represent distinct social phenomena, that share a fundamental characteristic: they can all be represented as temporal interactions involving two or more participants. This relational perspective lies at the heart of social network analysis and distinguishes it from other analytical frameworks (Wasserman, 1994). In an event-based representation of temporal networks, social dynamics are modeled as collections of time-stamped edges that capture the evolving structure of relationships over time (Lambiotte and Masuda, 2016). The aforementioned examples often involve relationships that extend beyond simple dyadic interactions, consisting of complex and multi-actor activities occurring at specific points in time. We refer to these richer forms of interaction as hyperevents (Lerner et al., 2021) — mathematically represented as time-stamped hyperedges (Berge, 1989), i.e., sets of nodes of arbitrary size. Modeling these hyperevents involves accounting for the polyadic nature of such interactions (Seidman, 1981).
@@
An increasing number of studies have focused on describing temporal hypergraphs using temporal motifs and relative counts (Paranjape et al., 2017; Wang et al., 2020; Lee and Shin, 2023). However, such approaches do not model the dynamics of how hyperedges emerge. Some progress has been made in detecting changes in the dynamics of temporal hypergraphs driven by specific factors (Zou and and, 2017). Notably, Benson et al. (2018) propose a generative stochastic model that predicts and explains the occurrence of sets of entities — interpretable as hyperevents — based on previously observed subsets. This growing interest in modeling complex, time-stamped group interactions has led to the extension of traditional Relational Event Models (REMs), which primarily focus on sequences of events involving a sender and a receiver (Butts, 2008; Bianchi et al., 2024), into the hypergraph domain. Relational Hyper Event Models (RHEMs) aim to capture the complex dynamics of polyadic events, conceived as interactions involving multiple participants simultaneously (Lerner et al., 2021; Lerner and Lomi, 2023). Going beyond the initial proposal by Perry and Wolfe (2013) for modeling multi-cast interactions, RHEMs provide a comprehensive framework for understanding the dynamics of social-relational hyper-event processes. Related intensity processes are modeled as functions of higher-order covariates (Lerner and Lomi, 2023). Recent RHEM formulations allow for the analysis of dependencies among nodes across different network modes, accommodating entities with inherently distinct characteristics — such as authors (scientists) and references (scientific papers) within a publication network (Lerner et al., 2025). The computational efficiency in calculating hyperedge covariates, enabled by the open-source software eventnet (Lerner and Lomi, 2020, 2023), makes RHEMs a practical tool in social network analysis. This facilitates the investigation of dynamic network effects — such as preferential attachment, (partial) repetition, and triadic closure — that may hold theoretical or empirical significance.
@@
Despite significant advancements in the RHEM literature, current formulations remain constrained by a fundamental limitation: they assume linearity in hyperevent effects. This reliance introduces two key drawbacks. First, it overlooks the temporal variability of effects. In temporal networks — where endogenous and complex dynamics are inherently time-dependent — several REM formulations have underscored the importance of incorporating time-varying effects into empirical applications (Bianchi and Lomi, 2023; Juozaitienė et al., 2023; Boschi et al., 2025). Second, the linearity assumption restricts the flexibility of existing REM frameworks, limiting their capacity to capture how covariate effects vary as a function of their values (Bauer et al., 2022; Filippi-Mazzola and Wit, 2024a; Filippi-Mazzola et al., 2024). This limitation is particularly relevant when covariates are evaluated in terms of their internal time (Juozaitienė and Wit, 2024; Amati et al., 2024), a common scenario for many endogenous covariates in relational event modeling. Mechanisms such as reciprocity and repetition — computed from previous events in the opposite and same direction respectively — can be more effectively assessed by considering the time intervals between current and past relevant events. While these prior interactions influence the rate of events, their impact typically weakens with time — a phenomenon referred to as forgetfulness (Juozaitienė and Wit, 2024). To capture this decaying influence, it is advantageous to model these covariates, measured as internal times, using non-linear effects. This allows the decay patterns to be learned directly from the data, rather than imposing a rigid, prespecified functional form. Such flexibility is helpful, because relying on an incorrect decay function can lead to erroneous conclusions (Arena et al., 2023). A related and equally important phenomenon is saturation, which arises when individuals face cognitive overload and become unable to effectively process additional information (Atienza-Barthelemy et al., 2025). Within the RHEM framework, saturation can be modeled by allowing the influence of certain drives on the event rate to exhibit a non-linear pattern — initially increasing with exposure but eventually reaching a plateau or declining.
@@
This paper addresses the primary limitation of current RHEM formulations by incorporating recent advancements from the REM literature, particularly regarding time-varying effects (TVE) and non-linear effects (NLE). However, our contribution goes beyond a straightforward adaptation of TVE and NLE to the context of relational hyper event modeling. We introduce a novel framework that enables effects to be jointly time-varying and non-linear. This is accomplished through the use of tensor product bases, which allow for the construction of smooth functions involving multiple covariates simultaneously (Wood, 2017). As a result, the proposed method supports the representation of a smooth interaction term, defined on the Cartesian product of time and the covariate of interest.
@@
The primary contribution of this paper is presented in Section 2, which outlines both the modeling framework and the inference techniques for relational hyperevent sequences. Section 3 describes a simulation study using synthetic relational hyperevent data. The results show that when the true data-generating process involves time-varying effects, non-linear effects, or the two of them jointly, the linear model fails to capture these dynamics. In contrast, the proposed model successfully recovers these effects. Moreover, when the assumption of linearity holds, the proposed framework naturally reduces to the correct linear specification, recovering it as a special case. Section 4 presents empirical results on scientific collaboration, extending the models developed by Lerner et al. (2025) [^2] The code required to implement the empirical analysis is provided in the Supplementary Material.. In this application, we find evidence for effects that are linear and fixed in time, non-linear, and jointly time-varying and non-linear, demonstrating the relevance of this approach in empirical hyperevent applications that take place over a long period — 80 years in our case.
@@ 2 Dynamic Hypernetwork Modeling
where $t_{m}$ is the time when a group of senders $I_{m}$ interacts with a group of receivers $J_{m}$. All participants are nodes in a vertex set $V$ of a temporal hypergraph. Depending on the context, senders and receivers may belong to the same or different groups. If they belong to the same group, the system is a one-mode system. If they play different roles, the system is a two-mode system. In the two-mode case, the vertex set $V$ of the two-model temporal hypergraph is divided into two disjoint sets: $V^{I}$ for possible senders and $V^{J}$ for possible receivers. If entities join or leave the system over time, these sets can vary with time and are written as $V^{I}_{t}$ and $V^{J}_{t}$. In our empirical study, we focus on a two-mode system related to scientific publications, where a group of authors (senders) cite a group of previously published papers (receivers). The notation $(t_{m},I_{m},J_{m})$ in Equation 1 naturally describes directed hyperevents, where sender and receiver groups are distinct. However, some interactions do not involve such roles. For example, in a meeting, participants interact together at the same time without a sender-receiver distinction. In such cases, we can describe the event as an undirected hyperevent by setting $J_{m}=\emptyset$ and including all participants in $I_{m}$. We denote the maximum size allowed for a meeting as $w$. Mathematically, undirected hyperevents represent undirected hyperedges, which are subsets of the vertex set $V$, while directed hyperevents correspond to directed hyperedges.
@@ 2.1 Relational Hyperevent Model
The event sequence in Equation (1) is treated as a realization of a marked point process. In this process, the points represent the times $t$ when hyperevents occur, and the marks specify the senders $I$ and receivers $J$ involved in each event. Related to this point process is a counting process $N(t,I,J)$, which records the number of hyperevents from group $I$ to group $J$ observed up to time $t$. The dynamics of the counting process can be modeled using both exogenous and endogenous features. Here, exogenous refers to factors external to the temporal hypergraph, while endogenous refers to factors arising from within it. To capture these dynamics, we incorporate the event history at time $t$ in the sub-$\sigma$-algebra $\mathcal{H}_{t}=\sigma\left(\{(t_{m},I_{m},J_{m})\mid t_{m}\leq t\}\right)$. If exogenous information is available, it should be included in the filtration $\mathbb{H}=\{\mathcal{H}_{t}\}_{t\geq 0}$.
Under the assumptions that make $N(\cdot,I,J)$ a right-continuous sub-martingale, adapted to the filtration $\mathbb{H}$, we can apply the Doob-Meyer theorem (Perry and Wolfe, 2013) to decompose it into two parts: a noisy component $M(\cdot,I,J)$, and a predictable component, the cumulative intensity process $\Lambda(\cdot,I,J)$. If it exists, the intensity function is defined as $\lambda(t,I,J)=\dfrac{\partial\Lambda(t,I,J)}{\partial t}$. This function describes the instantaneous rate at which the hyperevent $(t,I,J)$ occurs. We decompose this rate as follows:
@@
where $\mathcal{H}_{t^{-}}=\sigma\left(\{(t_{m},I_{m},J_{m})\mid t_{m}<t\}\right)$. This formulation is consistent with existing relational event and hyperevent models (Bianchi et al., 2024). The term $W(t,I,J)$ is the risk indicator, which determines whether the pair $(I,J)$ is at risk of a hyperevent at time $t$. The term $\lambda_{0}(t)$ is the baseline hazard, representing the baseline rate of event occurrence, which may vary over time. The main focus of this paper is the edge-specific contribution function, $f\left(\bm{x}(t,I,J),t\right)$, which captures the effect of covariates on the event rate. In the next section, we discuss its specification and interpretation in detail.

### v3

@@ Abstract
List of Abbreviations: REM: Relational Event Model RHEM: Relational Hyper Event Model DyNAM: Dynamic Network Actor Model TVE: Time-Varying Effect NLE: Non-Linear Effect ERGM: Exponential Random Graph Model LE: Linear Effect TVNLE: Jointly Time-Varying Non-Linear Effect AAS: Average Author Score RPI: Referenced Papers’ Impact SCAM: Shape Constrained Additive Models TPRS: Thin Plate Regression Spline TPS: Thin Plate Spline GLM: Generalized Linear Model GAM: Generalized Additive Model ECDF: Empirical Cumulative Distribution Function KS: Kolmogorov–Smirnov AIC: Akaike Information Criterion LogLik: Log-Likelihood
@@ 1 Introduction
Activities such as sending e-mails (perry2013point; boschi2024goodness), executing financial transactions (bianchi2023ties), attending meetings (lerner2021dynamic), citing academic references (filippi2024stochastic; lerner2024relational), co-participating in criminal activities (bright2024examining), and collaborating in cultural production (burgdorf2024communities) represent distinct social phenomena, that share a fundamental characteristic: they can all be represented as temporal interactions involving two or more participants. This relational perspective lies at the heart of social network analysis and distinguishes it from other analytical frameworks (wasserman1994social). In an event-based representation of temporal networks, social dynamics are modeled as collections of time-stamped edges that capture the evolving structure of relationships over time (lambiotte2016guide). The aforementioned examples often involve relationships that extend beyond simple dyadic interactions, consisting of complex and multi-actor activities occurring at specific points in time. We refer to these richer forms of interaction as hyperevents (lerner2021dynamic) – mathematically represented as time-stamped hyperedges (berge1989hypergraphs), i.e., sets of nodes of arbitrary size. Modeling these hyperevents involves accounting for the polyadic nature of such interactions (seidman1981structures).
@@
An increasing number of studies have focused on describing temporal hypergraphs using temporal motifs and relative counts (paranjape2017motifs; wang2020efficient; lee2023temporal). However, such approaches do not model the dynamics of how hyperedges emerge. Some progress has been made in detecting changes in the dynamics of temporal hypergraphs driven by specific factors (Zou2017Modeling). Notably, benson2018sequences propose a generative stochastic model that predicts and explains the occurrence of sets of entities – interpretable as hyperevents – based on previously observed subsets. This growing interest in modeling complex, time-stamped group interactions has led to the extension of traditional Relational Event Models (REMs), which primarily focus on sequences of events involving a sender and a receiver (butts2008relational; bianchi2024relational), into the hypergraph domain. Relational Hyper Event Models (RHEMs) aim to capture the complex dynamics of polyadic events, conceived as interactions involving multiple participants simultaneously (lerner2021dynamic; lerner2023relational). Going beyond the initial proposal by perry2013point for modeling multi-cast interactions, RHEMs provide a comprehensive framework for understanding the dynamics of social-relational hyper-event processes. This modeling approach allows us to address several important empirical questions. Consider co-offending, which can be treated as a relational hyper-event because it involves multiple individuals. Relational hyper-event models can explain co-offending dynamics as a function of the age and sex of the individuals involved, homophily with respect to these characteristics, and prior solo offending activity (bright2024examining). The same idea extends naturally to other social phenomena discussed at the beginning of this introduction. For example: How do reciprocity and transitive closure influence email exchanges? How does transaction amount or country of the beneficiary account influence the rate of bank transfers? Such questions can be addressed by modeling the intensity of the corresponding social processes as a function of observed covariates.
In relational hyper event modeling, intensity processes are indeed modeled as functions of higher-order covariates (lerner2023relational). Recent RHEM formulations allow for the analysis of dependencies among nodes across different network modes, accommodating entities with inherently distinct characteristics – such as authors (scientists) and references (scientific papers) within a publication network (lerner2024relational). The computational efficiency in calculating hyperedge covariates, enabled by the open-source software eventnet (lerner2020reliability; lerner2023relational), makes RHEMs a practical tool in social network analysis. This facilitates the investigation of dynamic network effects – such as preferential attachment, (partial) repetition, and triadic closure – that may hold theoretical or empirical significance.
Despite significant advancements in the RHEM literature, current formulations remain constrained by a fundamental limitation: they assume linearity in hyperevent effects. This reliance introduces two key drawbacks. First, it overlooks the temporal variability of effects. In temporal networks – where endogenous and complex dynamics are inherently time-dependent – several REM formulations have underscored the importance of incorporating time-varying effects into empirical applications (juozaitiene2023analysing; boschi2023smooth). The work of kamalabad2023point further motivates the need to account for changes in the impact of covariates in REMs. This need arises due to potential changes in the social interaction behavior of the entities in a temporal network. They test specific time points at which such changes in the parameters in the model formulation may be necessary. Within the broader framework of modeling relational event data, time-varying effects have also been incorporated into Dynamic Network Actor Models (DyNAMs) (uzaheta2024modeling). The underlying motivation is that the temporal network may occupy multiple states over time, and that the influence of network features on actor choices may vary across these states.
@@
Second, the linearity assumption restricts the flexibility of existing REM frameworks, limiting their capacity to capture how covariate effects vary as a function of their values (bauer2022smooth; filippi2024modeling; filippi2024analyzing). This limitation is particularly relevant when covariates are evaluated in terms of their internal time (juozaitiene2024nodal; amati2024goodness), a common scenario for many endogenous covariates in relational event modeling. Mechanisms such as reciprocity and repetition – computed from previous events in the opposite and same direction, respectively – can be more effectively assessed by considering the time intervals between current and past relevant events. While these prior interactions influence the rate of events, their impact typically weakens with time – a phenomenon referred to as forgetfulness (juozaitiene2024time). To capture this decaying influence, it is advantageous to model these covariates, measured as internal times, using non-linear effects. This allows the decay patterns to be learned directly from the data, rather than imposing a rigid, prespecified functional form. Such flexibility is helpful because relying on an incorrect decay function can lead to erroneous conclusions (arena2023fast). A related and equally important phenomenon is saturation, which arises when individuals face cognitive overload and become unable to effectively process additional information (atienza2025modeling). Within the RHEM framework, saturation can be modeled by allowing the influence of certain drivers on the event rate to exhibit a non-linear pattern – initially increasing with exposure but eventually reaching a plateau or declining. Indeed, we further emphasize that, for a linear model, the issue is not only non-linearity but especially non-monotonicity. A linear function can oversimplify a monotonic non-linear trend while still capturing its overall direction. However, it is inherently unable to represent a non-monotonic pattern. In some cases, for example, over-exposure may even reverse the direction, making it counter-effective. A linear model would be unable to reflect such a reversal and would instead provide a single estimate that “averages” the increasing and decreasing tendencies observed in the data.
@@
This paper addresses the primary limitation of current RHEM formulations by incorporating recent advancements from the REM literature, particularly regarding time-varying effects (TVEs) and non-linear effects (NLEs). However, our contribution goes beyond a straightforward adaptation of TVEs and NLEs to the context of relational hyper event modeling. We introduce a novel framework that enables effects to be jointly time-varying and non-linear. This is accomplished through the use of tensor product bases, which allow for the construction of smooth functions involving multiple covariates simultaneously (wood2017generalized). As a result, the proposed method supports the representation of a smooth interaction term, defined on the Cartesian product of time and the covariate of interest.
@@
The rest of the paper is organized as follows: We expand the motivation for this contribution in Section 2. The methodological section of the paper is presented in Section 3, which outlines the modeling framework and inference techniques for relational hyperevent sequences. Section 4 describes a simulation study using synthetic relational hyperevent data. The results show that when the true data-generating process involves time-varying effects, non-linear effects, or the two of them jointly, the linear model fails to capture these dynamics. In contrast, the proposed model successfully recovers these effects. Moreover, when the assumption of linearity holds, the proposed framework naturally reduces to the correct linear specification, recovering it as a special case and without overfitting. Section 5 presents empirical results on scientific collaboration, extending the models developed by lerner2024relational. In this application [^2] The code required to implement the empirical analysis is provided in the Supplementary Material., we find evidence for effects that are linear and fixed in time, but also – and especially – non-linear, and jointly time-varying non-linear, demonstrating the relevance of this approach in empirical hyperevent applications that take place over a long period – 80 years in our case. For instance, we identify non-monotonic effects – such as for author self-citation – whose influence on publication rates changes over time and in some time-slots end up being clearly non-monotonic. Such patterns cannot be captured by linear specifications and motivate extending relational hyper event models beyond linear formulations.
2 Motivation
Science and technology are key components in shaping the direction of innovation (wang2021probing). Innovation is a complex phenomenon, and defining proper measures of innovation has long been a central interest among researchers (todorov2024economies). Among the earliest measures of technological innovation has been the simple count of patents (griliches1998patent), although this alone fails to capture the interconnectedness of innovation (jaffe2002patents). More recent approaches focus on patent citations, based on the idea that if a patent has been published, it likely builds on prior patented work referenced in the cited patents (filippi2024stochastic), thereby identifying a “knowledge flow.” Similarly, scientific innovation has been described as a process where citations tracks the diffusion of innovation (zhai2018measuring). Although citation networks have been described as static network, e.g., using Exponential Random Graph Models (ERGMSs) (chakraborty2020patent), the dynamic flow of citations is better captured using dynamic network modeling techniques such as REMs (artico2023dynamic).
Given the complex nature of the citation process, the need of moving beyond a linear REM was recognized early on (bauer2022smooth; filippi2024analyzing). This allows the discovery of interesting non-monotonic patterns that linear models would fail to capture. For example, filippi2024stochastic, based on an analysis of United States Patent and Trademark Office patent citation data, detect that the citation rate of an individual patent shows a clear peak appearing shortly after the year 2000, in contrast to previously reported steady increases in the patent citation process. Additionally, the temporal lag at which future patents are most likely to be cited also exhibits a peak, occurring around five years.
golosovsky2019citation proposes a Hawkes self-exciting, non-linear process to model the latent citation rate of a paper over time. They demonstrate the non-linearity of citation dynamics and justify it as closely related to network topology. In particular, the effect of a paper’s citation history on the rate of new citations must itself depend non-linearly on its past, reflecting the very different trajectories of moderately versus highly cited papers. singh2023forgotten consider another phenomenon called citational amnesia, defined as the tendency to “forget” to cite high-quality older work. Citational amnesia affects future citations in two distinct ways. First, the recency of past publications affects how likely they are to be cited. Second, acknowledging that the impact of historical publication activity may itself change as the scientific ecosystem evolves is essential. Allowing for both non-linearity (to capture non-linear effects of past activity) and time-variation (to incorporate the historical drivers of citation amnesia) is therefore crucial.
Non-linear effects in relational dynamics happens in various situations. (juozaitiene2024time) show the importance of non-linear reciprocity and transitive closure effects in various communication networks. (boschi2024goodness) examine email communication data in a manufacturing company, demonstrating that the tendency to reply to an email depend in a strongly non-linear way on the time since the last email received from that individual. An example on time-varying effects in relational event models can be found in the effect of trade on alien species diffusion process over the past 150 years (juozaitiene2023analysing; boschi2023smooth). The effect of trade in this context is shown to vary over time: the diffusive effect of trade has gradually been decreasing over time. A possible explanation is that the nature of traded goods has changed, becoming less closely relevant for the invasion processes in recent years. The work of kamalabad2023point further underlines the need for testing the presence of change points and the consequent modification of the model formulation to account for time-varying effect. uzaheta2024modeling extend a special class of dynamic network models, DyNAMs, to allow for time-varying parameters, thereby accounting for changes in the system over time. Specifically, they assume that the state of the network may change over time, and depending on the state, the network features may affect the rate and choice sub-models that constitute a DyNAM differently. They apply their method to evaluate communication patterns in a MIT Media Lab’s “Social Evolution” study.
@@ 3 Dynamic Hypernetwork Modeling
[…]
Figure 1: Motivating example: publication events. This picture represent the first three events of a synthetic relational hyperevent sequence of paper publications. Each directed hyperevent consists of a paper, written by a team of authors (senders) and citing a set of references (receivers). Each hyperevent comes with a time-stamp, reporting the publication date. Only when the paper is published, it can be cited. This is why published paper appears as potential paper in the following time stamp. Furthermore, each actor is associated with an score (S), while each paper has a numeric impact associated (I).
@@
where $t_{m}$ is the time when a group of senders $I_{m}$ interacts with a group of receivers $J_{m}$. All participants are nodes in a vertex set $V$ of a temporal hypergraph. Depending on the context, senders and receivers may belong to the same or different groups. If they belong to the same group, the system is a one-mode system. If they play different roles, the system is a two-mode system. In the two-mode case, the vertex set $V$ of the two-model temporal hypergraph is divided into two disjoint sets: $V^{I}$ for possible senders and $V^{J}$ for possible receivers. If entities join or leave the system over time, these sets can vary with time and are written as $V^{I}_{t}$ and $V^{J}_{t}$. In our empirical study, we focus on a two-mode system related to scientific publications, where a group of authors (senders) cite a group of previously published papers (receivers).
Throughout this paper, we will refer to a simple synthetic example to introduce our methodological contributions from a practical point of view, relying on the scientific empirical application presented later. Figure 1 reports the first three publication events in a synthetic event sequence $E$ spanning from 1950 to 2021. At each time point, we observe a set $V^{I}$ of potential senders and a set $V^{J}$ of potential receivers. Specifically, we consider a set of hypothetical hyperevents as shown also in Table 1. This example is useful for highlighting the importance of defining a proper time-evolving risk set: indeed, a paper cannot cite another paper that has not yet been published.
Index $m$ Time $t_{m}$ Potential Authors $V^{I}_{t_{m}}$ Observed Authors $I_{m}$ Article $j_{m}$ Potential Citations $V^{J}_{t_{m}}$ Observed Citations $J_{m}$ 1 1950 $\{1,2,3,4\}$ $\{2,3\}$ $C$ $\{A,B\}$ $\{A,B\}$ 2 1951 $\{1,2,3,4\}$ $\{1,2,4\}$ $D$ $\{A,B,C\}$ $\{A,C\}$ 3 1953 $\{1,2,3,4\}$ $\{4\}$ $E$ $\{A,B,C,D\}$ $\{A,C\}$
Table 1: Collection of hyperevents $\{(t_{m},I_{m},J_{m})~|~m=1,\ldots,2000\}$ representing published papers whereby a set of co-authors $I_{m}$ cites a set of papers $J_{m}$ in a published article $j_{m}$.
The notation $(t_{m},I_{m},J_{m})$ in Equation 1 naturally describes directed hyperevents, where sender and receiver groups are distinct. The three directed hyperevents described in Table 1 can be represented in a two-mode network visualization. As shown in Figure 1, each event is a time-stamped directed hyperedge, originating from a set of nodes of one type and pointing to a set of nodes of another type. However, since we have sets of authors linked to sets of cited papers in a single event, we also get relations between authors (co-authorship) and between papers (being co-cited). This allows for the computation of endogenous covariates that are functions of previously occurred relations of the type: (i) paper-cites-paper, (ii) author-cites-paper, (iii) author-cites-author, (iv) author-publishes-paper.
There also exist relational hyperevents with a sender-receiver sets structure. For example, in a meeting, participants interact together at the same time without a sender-receiver distinction. In such cases, we can describe the event as an undirected hyperevent by setting $J_{m}=\emptyset$ and including all participants in $I_{m}$. We denote the maximum size allowed for a meeting as $w$. Mathematically, undirected hyperevents represent undirected hyperedges, which are subsets of the vertex set $V$.
@@ 3.1 Relational Hyperevent Model
The event sequence in Equation (1) is treated as a realization of a marked point process. In this process, the points represent the times $t$ when hyperevents occur, and the marks specify the senders $I$ and receivers $J$ involved in each event. Related to this point process is a counting process $N(t,I,J)$, which records the number of hyperevents from group $I$ to group $J$ observed up to time $t$. The dynamics of the counting process can be modeled using both exogenous and endogenous features. Here, exogenous refers to factors external to the temporal hypergraph, while endogenous refers to factors arising from within it. To capture these dynamics, we incorporate the event history at time $t$ in the sub-$\sigma$-algebra $\mathcal{H}_{t}=\sigma\left(\{(t_{m},I_{m},J_{m})\mid t_{m}\leq t\}\right)$. If exogenous information is available – such as author score and paper impact in the synthetic example – it should be included in the filtration $\mathbb{H}=\{\mathcal{H}_{t}\}_{t\geq 0}$.
Under the assumptions that make $N(\cdot,I,J)$ a right-continuous sub-martingale, adapted to the filtration $\mathbb{H}$, we can apply the Doob-Meyer theorem (perry2013point) to decompose the counting process into two parts: a noisy component $M(\cdot,I,J)$, and a predictable component, the cumulative intensity process $\Lambda(\cdot,I,J)$. If it exists, the intensity function is defined as $\lambda(t,I,J)=\dfrac{\partial\Lambda(t,I,J)}{\partial t}$. This function describes the instantaneous rate at which the hyperevent $(t,I,J)$ occurs. We decompose this rate as follows:
@@
where $\mathcal{H}_{t^{-}}=\sigma\left(\{(t_{m},I_{m},J_{m})\mid t_{m}<t\}\right)$. In this paper, we use both the term rate and the term hazard to indicate $\lambda(t,I,J;\mathcal{H}_{t^{-}})$. This formulation is consistent with existing relational event and hyperevent models (bianchi2024relational). The term $W(t,I,J)$ is the risk indicator, which determines whether the pair $(I,J)$ is at risk of a hyperevent at time $t$. The term $\lambda_{0}(t)$ is the baseline hazard, representing the baseline rate of event occurrence, which may vary over time. The main focus of this paper is the edge-specific contribution function, $f\left(\bm{x}(t,I,J),t\right)$, which captures the effect of covariates on the event rate. In the next section, we discuss its specification and interpretation in detail.